OpenStudy (anonymous):

can you please explain this fraction 4 1/3 divided 3 1/2 please

OpenStudy (anonymous):

4 1/3 as a fraction is 13/3. 3 1/2 as a fraction is 7/2. 13/3 divided by 7/2 is 13*2/(2*7) = 13/7.

OpenStudy (anonymous):

will you please draw it?

OpenStudy (anonymous):

Sorry no drawing, just read.

OpenStudy (anonymous):

okay Thank you !

OpenStudy (anonymous):

6x+22=3x+7

OpenStudy (anonymous):

13/7